OpenStudy (anonymous):

In the sentence below, which word or phrase refers to a mathematical operation? While driving on the freeway, Jennifer was operating her vehicle at 10 miles per hour more than Amy. A. more than B. miles C. 10 D. operating

OpenStudy (biokiiller396):

A because D refers to her using her car C refers to The amount of miles And B is just length and A means you will have to subtract

OpenStudy (anonymous):

Which equation represents two less than a number is equal to 10? A. n – 2 = 10 B. n + 2 = 10 C. 2 – n = 10 D. 2 + n = 10

OpenStudy (biokiiller396):

A because 2 will be subtracted FROM a number... Subtracted from = less than a number